OpenStudy (anonymous):

Please help I give medals >.< The index of refraction of a transparent substance is 1.7, and the speed of light in air is 3.00 · 108 s. Calculate the speed of light in the transparent substance.

OpenStudy (anonymous):

@ShailKumar

OpenStudy (anonymous):

@AkashdeepDeb @Abhisar

OpenStudy (abhisar):

Hello @Miguel27525 !

OpenStudy (anonymous):

Hello

OpenStudy (abhisar):

\(\huge n = \frac{c}{v}\)

OpenStudy (abhisar):

1.7=(3*10\(^8\))/v

OpenStudy (abhisar):

V=1.7*10\(^8\)

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

17000000 right?

OpenStudy (abhisar):

17 followed by 7 zeros, 170000000